# Towards Generalised Pre-Training of Graph Models

### v3

@@ Abstract
The principal benefit of unsupervised representation learning is that a pre-trained model can be fine-tuned where data or labels are scarce. Existing approaches for graph representation learning are domain specific, maintaining consistent node and edge features across the pre-training and target datasets. This has precluded transfer to multiple domains.
In this work we present Topology Only Pre-Training, a graph pre-training method based on node and edge feature exclusion. Separating graph learning into two stages, topology and features, we use contrastive learning to pre-train models over multiple domains. These models show positive transfer on evaluation datasets from multiple domains, including domains not present in pre-training data. On 75% of experiments, ToP models perform significantly ($P\leq 0.01$) better than a supervised baseline. These results include when node and edge features are used in evaluation, where performance is significantly better on 85.7% of tasks compared to single-domain or non-pre-trained models. We further show that out-of-domain topologies can produce more useful pre-training than in-domain. We show better transfer from non-molecule pre-training, compared to molecule pre-training, on 79% of molecular benchmarks.
@@ 1 Introduction
Much work has been conducted on forming useful representations of graphs which are more easily used by well-explored methods. Khoshraftar et al. [15], Hamilton et al. [12] and Chen et al. [4] survey graph representation learning in general. Such representation learning, after transfer and fine-tuning, allows greater performance in areas of data or label scarcity.
Whilst there is a good deal of research into graph representation learning methods, there is little work on methodologies to combine different domains of graph data effectively. Ubiquitous in other fields, primarily text and images, is pre-training on a large amalgamated dataset of samples, aiming to span the full or realistic space of the data [28]. The main obstacle to such an approach for graphs is consistency in node and edge features, as graph data presents a large diversity in possible feature-sets.
Given this variety of features between domains current research argues, but does not show, that multi-domain pre-training on graphs is not feasible [21, 31, 39]. There are very few works attempting learning across multiple domains of graphs [19]. This excludes areas of data or label-sparsity from the benefits of pre-training. Multi-domain pre-training will allow learning on domains where data is too scarse to form pre-training datasets.
Here we propose Topology Only Pre-Training (ToP), a generalised pre-training method for graph models, under which pre-training is conducted with node and edge features excluded. We create a multi-domain dataset, and using ToP through adversarial contrastive learning, produce pre-trained models with consistently positive transfer across multiple domains and tasks. Transfer is also positive when node and edge features are included in downstream datasets during fine-tuning.
@@
In this work we propose and investigate ToP as a pre-training method for graph data that is not limited to a given domain or task. The result of such pre-training is a single model that can be transferred with broadly positive performance increases on arbitrary graph tasks and domains. Experimentally we test two hypotheses:
@@
With the exclusion of node and edge features out-of-domain pre-training can have greater benefits than in-domain pre-training across task-relevant metrics.
Findings We evaluate ToP using graph constrastive learning with both random and adversarial augmentations. Through this evaluation we contribute the following empirical findings:
@@
    Effective pre-training without features. Pre-training on varied domains, under the exclusion of node and edge features, leads to positive transfer. We find that under fine-tuning, results are at worst on-par, and on 76% of tasks significantly better, than a non-pretrained baseline.
@@
    Increased performance from domain diversity. Pre-training without features from varied non-target domains, instead of the target domain, can bring greater downstream performance benefits. On molecular benchmarks, we find that a model pre-trained on non-molecules, compared to pre-trained only on molecules, is at worst on-par. On 56% of tasks the non-molecule model performs significantly better, and never significantly worse, than the molecule model.
Given these findings, under recent definitions, these may be among the first foundation graph models [2, 19, 21, 39]. Both of these findings go beyond the current consensus of graph pre-training, opening broad avenues of research and application. In particular our second finding runs contrary to the single-domain training paradigm.
@@
Organisation In Section 2, we describe the notation and concepts necessary to understand our method and provide an overview of related works. In Section 3, we describe our generalised pre-training method ToP. In Section 4, we describe how we evaluate the hypotheses detailed above, then detail the experiments we perform to evaluate ToP and these hypotheses. In Section 5 we discuss the implications of the results from these experiments, as well as list the assumptions we make and threats to validity. In Section 6, we conclude with avenues for future research.
2 Background and Related Works
@@
Here we describe the concepts necessary to understand the rest of this work.
@@
Representation Learning Let $X\in\mathcal{X}$ represent a random variable $X$ in the data space $\mathcal{X}$, and $Y\in\mathcal{Y}$ the same for the output space. Representation learning aims, given their joint distribution $P(X,Y)$, to learn a mapping $X\rightarrow\hat{X}$ that maximises their mutual information $I(X;Y)$. In most contexts this means producing numerical vectors for input data that usefully describe that data.
@@
Contrastive Learning Let $\mathcal{T}$ be set of small transformations called ‘augmentations’, where $t\in\mathcal{T}:X\rightarrow X^{\prime}$. These augmentations are generally designed to manipulate the input sample in a way non-destructive to the ‘useful’ information it represents. Common choices for images, for example, are gray-scaling and rotations. Given an encoder’s mapping $F(X_{i})\rightarrow\hat{X_{i}}$, multiple augmentations of the same input sample should be as similar as possible.
@@
Graph Representation Learning Hamilton [11] provides a comprehensive summary of methods for Graph Representation Learning (GRL). Traditional methods to encode graphs as vectors include graph statistics such as centrality measures and clustering coefficients, but also kernel methods such as the Weisfeiler-Lehman (WL) kernel [29]. Expressive Graph Neural Networks (GNNs) models, generally through parameterised message-passing, can produce powerful representations of graphs. A number of GNN architectures have been proposed with the Graph Isomorphism Network (GIN) being the most expressive [41] in representation learning contexts.
@@
Graph Contrastive Learning Multiple studies exist for graph contrastive learning (GCL) [5, 10, 13, 14, 18, 38, 42]. The typical graph contrastive learning formulation, Graph Contrastive Learning (GraphCL), is put forward by You et al. [38]. In GraphCL and many subsequent works, such augmentations are performed randomly using fixed parameters, i.e., ‘drop 20% of nodes’.
@@
Adversarial Graph Contrastive Learning It has been shown that the quality of augmentations is critical for graph contrastive learning [38, 42]. To maximise expressivity in these augmentations adversarial strategies have been proposed [13, 40]. Suresh et al. [32] propose Adversarial-Graph Contrastive Learning (AD-GCL), with their method training both an encoder and a ‘view-learner’ $\mathcal{T}_{\theta}$. AD-GCL performs the standard GCL protocol with weight updates to a parameterised ($\theta$) set of augmenting functions $\mathcal{T_{\theta}}$ through the contrastive loss. Here $\mathcal{T_{\theta}}$ learns augmentations that minimise mutual information between the source and augmented graphs through edge dropping. The encoder, adversarially to this, learns to maximise shared information between the respective augmentations. The updates for $\mathcal{T}$ are regularised such that augmentations are not entirely destructive.
One For All A work concurrent with ours is One-For-All (OFA) [19]. OFA employs a single model on graph and node classification tasks across domains, sidestepping the issues around node and edge features we’ve discussed by assuming that natural language descriptions of features are adequate, and using language model embeddings as features. OFA, then, requires that input graphs have textual features, or that existing features can be usefully described in natural language. Whereas OFA produces only classification models, ToP produces pre-trained models intended for downstream fine-tuning. As such ToP allows arbitrary downstream tasks to be conducted which are not possible with OFA, regression being an example.
@@ 3 ToP: Topology Only Pre-Training
Figure 1: The process for handling features with ToP. During pre-training, features are replaced with a single integer label for each node and edge (all identical), which are then passed to a single-layer MLP. This results in an identical input vector for each node. During transfer, a different input head is used, moving the original dimensions to the hidden dimensionality of the GNN encoder block. This allows arbitrary node and edge features to be included in transfer. Omitting the output head of the model allows transfer onto node and edge-level tasks.
One naive approach to generalised pre-training with varied feature sets included would be to pass features as-is, with padded vectors where feature sets are of a lower dimensionality. Here components would be shared across domain features, for example types of atoms and number of social followers in the same component, representing an extremely difficult space to learn. Alternatively one could reserve components of a long feature vector for each feature set, with padding tokens in components for other domains and feature sets. This would require a model to learn a massive span of features, and preclude movement to new feature sets, and again learning is unlikely to occur. Feature-management approaches, which move domain features to the same feature set [19], are in their infancy.
@@
Given that pre-training while including features in their original form is not feasible, we view graph learning as a task occurring on both topologies $G$ and features on that topology $X$. From this view, pre-training without features is well-motivated, with models learning generalised information about graph structures before features are introduced. This information from $G$ is useful for downstream tasks, even while missing task-specific information, and in situations where data is abundant pre-training on both $X$ and $G$ will likely still lead to higher performance.
This is an abstraction away from feature-specific correlations, with abstraction a core quality of effective representation learners [16]. Pre-training without topologies, viewed as an initialisation, would provide a smoother loss landscape for downstream tasks [20]. In this case a wider range of topologies could represent more useful information than a narrow range closer to a downstream domain.
@@
Topology Only Pre-Training (ToP) is a method for generalised pre-training of graph models. Under ToP pre-training is conducted only on topologies, with node and edge features excluded or replaced with identical values. This use of only topologies $G$, instead of graphs with specific features $\underline{G}$, allows pre-training on a multi-domain dataset. Prior to this work, the use of such a graph dataset with mixed node and edge features was not considered feasible [21, 31, 39].

### v4

@@ Abstract
The principal benefit of unsupervised representation learning is that a pre-trained model can be fine-tuned where data or labels are scarce. Existing approaches for graph representation learning are domain specific, maintaining consistent node and edge features across the pre-training and target datasets. This has precluded transfer to multiple domains. We present Topology Only Pre-Training (ToP), a graph pre-training method based on node and edge feature exclusion. We show positive transfer on evaluation datasets from multiple domains, including domains not present in pre-training data, running directly contrary to assumptions made in contemporary works. On 75% of experiments, ToP models perform significantly ($P\leq 0.01$) better than a supervised baseline. Performance is significantly positive on 85.7% of tasks when node and edge features are used in fine-tuning. We further show that out-of-domain topologies can produce more useful pre-training than in-domain. Under ToP we show better transfer from non-molecule pre-training, compared to molecule pre-training, on 79% of molecular benchmarks. Against the limited set of other generalist graph models ToP performs strongly, including against models with many orders of magnitude larger. These findings show that ToP opens broad areas of research in both transfer learning on scarcely populated graph domains and in graph foundation models.
@@ 1 Introduction
Representation learning allows models to convert the useful information from rich data, for example text and images, into flexible vector encodings. As pre-training, then fine-tuning on downstream tasks, representation learning leads to improved performance [6, 22, 9]. This is an invaluable technique where downstream labels are scarce, allowing expressive models to be applied where otherwise performance would be unsatisfactory. Graph representation learning extends these methods to structured data, allowing much work on unsupervised learning over large datasets of graphs from a given domain [16, 5, 21].
In other domains, such as images and text, it is common to pre-train on a large amalgamated dataset of samples, aiming to maximise coverage of the space [38, 45, 33]. Large models pre-trained on these massive datasets then show strong performance on a large range of downstream tasks. Models with this combination of large size, large data and strong performance are commonly referred to as foundation models [30].
There are substantial obstacles to taking this approach for graph data, the largest being feature heterogeneity between domains. Images or text from multiple domains are fundamentally the same data-type, uniform-size arrays of pixel values or sequences of tokens respectively, meaning that the multi-domain problem can be tackled head-on through data and model scale. Graphs, on the other hand, have a large diversity of possible feature sets. We present model examples of feature sets for social networks and molecules in Figure 1. How a single model can handle highly heterogenous feature sets like these, for example both a text biography and the atomic number of an atom, is an open and challenging problem.
Figure 1: Graph data has rich, highly varied feature sets between domains. Here we provide model examples from social media (left) and chemistry (right). Social media feature sets might include a combination of textual, image and numeric types. Chemical feature sets are primarily numeric and categorical, representing properties of component atoms that might be difficult to infer for deep-learning models.
This has lead current research to argue or assume, but not show, that multi-domain pre-training on graphs is not feasible [26, 42, 56]. As a consequence there are few works attempting learning across multiple domains of graphs. Some very recent works have proposed approaches using LLMs as feature transformations [40, 55, 52, 25], effectively moving a source domain into a unified text-featured domain. This drastically increases computational expense, and when features are not easily or commonly expressible as text, for example continuous features, LLM approaches presumably cannot be used.
In downstream tasks feature information is often essential, so generalised approaches should allow them to be included in downstream applications, unlike in the very limited set of approaches for generalised graph pre-training [32]. The lack of a generalised pre-training method, which allows domain features to be included downstream, excludes areas of data sparsity or scarcity from the benefits of pre-training. Even on areas where data is plentiful enough for pre-training, it requires a different pre-trained model to be used each time, increasing the time, complexity, and cost of usage. Multi-domain pre-training would avoid these issues, both opening transfer learning as an option for domains with limited access to data, and providing a baseline model on which applications can be built.
1.1 Pre-Training without Features
Graph learning can rely on information from both structure and features, with some of that information coming only from structure. Therefore, beneficial learning during pre-training can occur with features excluded. Motivated by this view we introduce Topology Only Pre-Training (ToP), a generalised pre-training method for graph models. Under ToP pre-training is conducted with node and edge features excluded, but during downstream application they can be re-included by swapping components of a given model.
@@ 1.1 Pre-Training without Features
We investigate ToP as a pre-training method for graph data that is not limited to a given domain or task. The result of such pre-training is a single model that can be transferred with positive performance increases on arbitrary graph-level tasks and domains. Experimentally we introduce and test two hypotheses:
@@
With the exclusion of node and edge features out-of-domain pre-training can have greater benefits than in-domain pre-training across downstream tasks.
Findings
We evaluate ToP using graph constrastive learning with both random and adversarial augmentations. This evaluation includes the construction of an amalgamated, multi-domain dataset, spanning from road networks to molecules. We pre-train multiple models on different subsets of this amalgamated dataset and with varied augmentation schemes. We perform multiple fine-tuning runs for each model in order to show statistical significance for our experiments. Through this evaluation we contribute the following empirical findings:
@@
    Effective pre-training. Pre-training on varied domains, under the exclusion of node and edge features, leads to positive transfer. We find that under fine-tuning, results are at worst on-par, and on 76% of tasks significantly better, than a non-pretrained baseline.
@@
    Increased performance from domain diversity. Pre-training without features from varied non-target domains, instead of the target domain, can bring greater downstream performance benefits. In the absence of features in pre-training a non-molecule learner consistently produces positive transfer compared to a molecule learner. On 56% of tasks the non-molecule model performs significantly better, and never significantly worse, than the molecule-only model.
Contributions
We present two main contributions, although the work contains several more.
Our first contribution is ToP, a multi-domain pre-training method for graph models, with original domain features included during downstream tasks. Therefore, and under recent definitions, models pre-trained using ToP are close to being foundation graph models [2, 25, 26, 56]. ToP is the first method that, to our knowledge, has the following desirable properties:
  1. 1.
    Positive transfer across multiple, highly diverse domains
  2. 2.
    Original node and edge features during downstream transfer
  3. 3.
    No very large auxiliary models
  4. 4.
    No downstream feature transformations
Across molecular benchmarks our implementation of ToP out-performs benchmarks from the limited set of generalist GNN approaches that exist. Despite not aiming to out-perform domain specific specialist models, ToP performs better-than or on-par with generalist LLMs many orders of magnitude larger in both parameters and training dataset size. This provides great utility when pre-training data is not abundant.
Our second contribution is a new research direction: with domain features excluded, diverse non-domain pre-training can be more valuable than in-domain. This runs contrary to the single-domain training paradigm, and directly opposite to the assumptions made in other works [28, 42]. In Mao et al [28]’s review of LLMs for for multi-domain graph models, enumerating why they show promise, it is assumed that “…training on [graphs from multiple domains] simultaneously shows no positive transfer benefit while increasing the risk of the negative transfer”. Our work shows that the opposite is true when pre-training is conducted only on structure. This implies that greater performance through ToP can be achieved simply by increasing the scale and diversity of the pre-training dataset.
Organisation
@@
In Section 2, we describe the notation and concepts necessary to understand our method and provide an overview of related works. In Section 3, we describe our generalised pre-training method ToP, and our hypotheses about it as a method. In Section 4, we describe how we evaluate the hypotheses detailed above, and present the findings of those experiments. In Section 5, we discuss the implications of the results from these experiments, as well as list the assumptions we make and threats to validity. In Section 6, we conclude with avenues for future research.
2 Background and Related Work
@@
Here we describe the concepts necessary to understand the rest of our method, as well as closely related works.
Representation Learning
@@
In most contexts, representation learning is the process of forming useful numerical encodings of non-tabular input data through an ‘encoder’ [1]. Let $X\in\mathcal{X}$ represent a random variable $X$ in the data space $\mathcal{X}$, and $Y\in\mathcal{Y}$ the same for the output space. Representation learning aims, given their joint distribution $P(X,Y)$, to learn a mapping $X\rightarrow\hat{X}$ that maximises their mutual information $I(\hat{X};Y)$. In basic terms, representation learners aim to transfer some piece of rich data, for example images, into a vector that describes that piece of data.
Contrastive Learning (CL)
@@
CL is a semi-supervised school of representation learning based on altering input samples [22]. Let $\mathcal{T}$ be a set of small transformations called ‘augmentations’, where $t\in\mathcal{T}:X\rightarrow X^{\prime}$. These augmentations are generally designed to manipulate the input sample in a way non-destructive to the ‘useful’ information it represents. Augmentations can be understood as a method to increase the semantic coverage in the input space. Common choices for images, for example, are gray-scaling and rotations. Intuitively, the augmentation defines an invariant transformation for the encoder to learn. Given an encoder’s mapping $F(X_{i})\rightarrow\hat{X_{i}}$, multiple augmentations of the same input sample should be as similar as possible. Again in simple terms, contrastive learners aim to represent slight variations on the same sample in similar ways, for example through cosine similarity.
Graph Representation Learning
@@
Hamilton [15] provides a comprehensive summary of methods for graph representation learning. Traditional methods to encode graphs as vectors include graph statistics such as centrality measures and clustering coefficients, but also kernel methods such as the Weisfeiler-Lehman kernel [39]. Expressive Graph Neural Networks (GNNs) models, generally through parameterised message-passing, can produce powerful representations of graphs. A number of GNN architectures have been proposed with the Graph Isomorphism Network (GIN) being the most expressive in message-passing representation learning contexts [59].
Graph Contrastive Learning
@@
Multiple studies exist for graph contrastive learning [8, 14, 17, 20, 24, 54, 60]. The typical graph contrastive learning formulation, Graph Contrastive Learning (GraphCL), is put forward by You et al [54]. In GraphCL and many subsequent works, such augmentations are performed randomly using fixed parameters, i.e., ‘drop 20% of nodes’. These augmentation strategies do not vary based on the graph-instance being augmented.
Adversarial Graph Contrastive Learning
@@
The quality of augmentations is critical for graph contrastive learning [54, 60]. To maximise expressivity in these augmentations adversarial strategies have been proposed [17, 57]. Suresh et al [43] propose Adversarial-Graph Contrastive Learning (AD-GCL), with their method training both an encoder and a ‘view-learner’ $\mathcal{T}_{\theta}$. AD-GCL performs the standard GCL protocol with weight updates to a parameterised ($\theta$) set of augmenting functions $\mathcal{T_{\theta}}$ through the contrastive loss. Here $\mathcal{T_{\theta}}$ learns augmentations that minimise mutual information between the source and augmented graphs through edge dropping. The encoder, adversarially to this, learns to maximise shared information between the respective augmentations. The updates for $\mathcal{T_{\theta}}$ are regularised such that augmentations are not entirely destructive.
Large Language Models on Graphs
Large Language Models (LLMs) are models pre-trained, typically generatively, across massive corpuses of natural language data. Relying on scaling through both dataset size and model size, such models are highly performant on natural language processing. The first foundation models, though trained in the same manner, were smaller and typically required fine-tuning for strong performance on a specific task. Bi-Directional Encoder Representations from Transformers (BERT) is the archetypal example here [9]. The most recent LLMs vary in size from a few billion parameters [10] to many hundreds of billions [33].
LLMs have in very recent research been applied to building multi-domain graph models, often termed Graph Foundation Models (GFMs) in their respective works. Some have argued that LLMs themselves, by default, may be GFMs [28]. LLMs carry the significant benefit of allowing one-shot or few-shot application on downstream tasks in some circumstances. However, given their size, LLMs can drastically increase the cost and complexity of using graph models.
The simplest approach for applying LLMs on graph data is simply to encode a graph structure in natural language, along with features. We model our description of this approach on Fatemi et al [12]. Here a graph structure is encoded via some schema, for example the edge $(Callum,Zack)$ might be encoded ‘Callum and Zack are connected’ in a social network. Fatemi et al [12] find that the choice of encoding schema, for example ‘is friends with’ vs ‘is connected to’, has a significant impact on performance, and that in-general this approach is limited.
Alternatively the performant capacity of LLMs as encoders has been used to unify the graph feature space across domains. By describing the features of a node or edge in natural language, the LLMs encodings of that new textual feature can be shared between domain feature sets [40, 55, 52, 25]. This assumes and requires that:
  1. 1.
    Original features are easily encoded in text
  2. 2.
    Natural language versions of features are meaningful
  3. 3.
    LLM encodings are in fact a coherent embedding space
For natural-language features, the approach is trivial, and numeric features are converted to text via a schema. An example for molecular graphs would be ‘An atom with X protons, Y covalent bonds, which (is/is not) part of a ring structure’. These textual features are then passed to an LLM to form numeric encodings as features. This has the significant advantage of allowing one-shot or few-shot downstream use, assuming that downstream text features are close to in-domain. The approach has yet to be applied to graphs with rich feature sets. We assume this because it is not clear how previous approaches would encode non-intuitive graph features textually, such as continuous ones. Many of these initial works use multiple domains where features are already text, and do not attempt to move other feature sets into textual features.
Another approach has emerged in LLMs for chemistry, where there are expressive sequential descriptions of molecular graphs commonly used in natural language [27, 58, 4]. Here LLM approaches use sequential SMILES encodings of molecules, for example O=C=O from CO2, to represent molecules at some point of their pipeline. SMILES strings, as a natural language expression of a graph structure, are a useful tool for LLM approaches. Most other graph structures, for example social networks, are much harder to express in natural language. A core assumption of these models, with strong evidence for validity through experimentation, is that forming datasets of text surrounding these SMILES encodings using chemical literature allows learning about chemical dynamics to occur. Where these models do not include explicit usage of graph structures, for example InstructMol [4] and ChemDFM [58], they require far larger models and datasets for the same performance.
ToP carries few of the downsides associated with LLMs for graph data, yet achieves many of the same goals. We show that ToP models have strong downstream performance across domains on structure-only tasks, and in comparison, LLMs perform far more weakly. Despite the suitability of molecular tasks for LLMs, as described above, ToP performs at a level comparable to the best-performing LLMs on molecular tasks. This comes at a massively smaller time and computational cost compared to LLMs. Our reasonably large GNN encoders complete pre-training in less than half a day, on consumer hardware, with fine-tuning and inference both feasible and fast on consumer laptops.
Graph Constrastive Coding
A similar work to ToP is Graph Contrastive Coding (GCC) [32]. GCC begins by generating subgraphs from various large graphs through ego sampling. As augmentations it used random walks with restarts, with embeddings of the seed node used for its contrastive loss, resulting in a local-focus model. This is in contrast to this work, where we focus on individual graph level representations, instead of pre-training as a route to tasks on large graphs.
As in this work, GCC’s primary aim is to sidestep the issue of varied feature sets between graph domains. ToP simply removes feature sets during pre-training. GCC goes a step further, replacing existing node and edge features with positional information, principally the graph Laplacian. While this has obvious merit in encoding structural information, models can learn from specific local-level features, allowing overfitting and other feature-space biases to occur. Our assumption, validated through experiment, is that such an approach also makes re-introduction of domain-specific features much less effective. Where we do not extend GCC by including domain features during transfer, using structural encodings as in the original work, performance on out-of-domain data is weak. On out-of-domain structure-only tasks, ToP models perform more strongly or on-par with GCC models, despite lacking the advantage of structural features included with GCC.
@@ 3 ToP: Topology Only Pre-Training
Assume a dataset with features $X\in\mathbf{R}^{|V|\times D}$ and similar for edges, topology as an edgelist $E:\{(v_{1},v_{2}),\ldots\}$, and graph labels $y$, with graphs $G:\{X,E\}$. $y$ is some target label. When a given model aims to learn some mapping $f(G)\rightarrow y$, in reality this is $f(X,E)\rightarrow y$. The usual graph learning assumption is that topology $E$ and features $X$ must be considered together for information to be useful. In other words their combination has mutual information with a target $y$, $I(y;X,E)$.
@@
We can assume that in the absence of features, structural information is still present, i.e. $I(y;E)\geq 0$. From this view, pre-training without features is well-motivated, with models learning generalised information about graph structures before features are introduced. This information from $E$ is useful for downstream tasks, even while missing task-specific information from features. This is an abstraction away from feature-specific correlations, with abstraction a core quality of effective representation learners [22]. Removing features $X$ necessarily removes their own information, but also the interdependence of features and structure:
$I(y;E,X)\geq I(y;X)+I(y;E)$ (1)
In situations where data is abundant, therefore, pre-training on both $X$ and $E$ will likely still to higher performance.
Graph without features are sampled from the same high dimensional space, which is highly constrained in real world graphs. As such they can be similar across highly varied domains, much as in the image space, even if they are entirely heterogenous with features included. On the local scale at which message passing GNNs operate, the same patterns may occur across any number of domains. In this case a wider range of topologies could span a wider range of the basic graph space, and so contain more useful information, than a narrow range closer to a downstream domain.
This pre-training on only topologies would give an initial bias towards structural features during downstream transfer. Such a structural bias is encouraged in other works, in particular those works on Topological GNNs [18, 7, 48]. Here layers or models themselves are designed to encourage learning from patterns in structure over patterns in features.
@@
Topology Only Pre-Training (ToP) is a method for generalised pre-training of graph models. Under ToP pre-training is conducted only on topologies, with node and edge features excluded through replacement with identical values. This use of only topologies $G$, instead of graphs with specific features $\underline{G}$, allows pre-training on a multi-domain dataset. In the works prior to ToP, the use of such a graph dataset was not tractable [26, 42, 56].
Figure 2 shows a schematic for ToP models. ToP’s implementation is elegant: Features are removed from nodes and edges and replaced with some arbitrary identical value. In our case all nodes and edges carry the integer value 1. With the input head some simple MLP, the choice of value has no significant impact, as the GNN layers are presented the same vector for all nodes and edges. An encoder is then pre-trained using some unsupervised representation learning method. In our implementation we use contrastive learning for pre-training, given its strong performance in representation learning on graphs.
By swapping the input head of the pre-trained model, arbitrary node and edge features can be included during transfer, allowing integration of feature information alongside structural during downstream transfer. In the same manner the output head of a model can be replaced or excluded entirely, allowing transfer onto node and edge-level tasks.
Figure 2: A schematic for ToP models. During pre-training, features are replaced with a single integer label for each node or edge (all identical), which are then passed to a single-layer MLP. This results in an identical input vector for each node and edge. During transfer, a different input head is used, moving the original dimensions to the hidden dimensionality of the GNN encoder block. This allows arbitrary node and edge features to be included in transfer. Omitting the output head of the model allows transfer onto node and edge-level tasks.